Minutes 060801

From GlueXWiki
Jump to: navigation, search

Attendees: Matt(Indiana), Elton, David, Simon (JLab), Dan Sober (CU), George, Blake, Rafael, Zisis (Regina)

  1. Business Arising
    • Action Items: The Action Items list was reviewed.
  2. Detector Setup/Mechanics
    • Indiana: IU, CMU and FSU machine shops are done (see Optical Assembly). Eric is away this week but the cart will be assembled and load tested. The light guide work is done. The mu-metal test will be done next week.
    • Regina: Crate is ready and packed and awaits paperwork from Shipping and Receiving. We will aim to have it at JLab a few days before Aug. 21 and it will be moved directly into Hall-B. The mu-metal shields will be shipped shortly from 'Mu Metal' to IU. George is working on the cosmics trigger counters.
  3. Hall-B Tasks:
    • Beamline/alcove: All work is proceeding as expected. We discussed obtaining black tedlar (or equivalent) to cover the module and PMTs against light. The IU light guide box has light baffles and black rubber gaskets and we plan to use black caulk around edges.
    • Cable and trigger timing: David will program the trigger supervisor for the cosmics triggers. The fast cable from the trigger MOR is expected next week. Comments/additions on the Beam Test Trigger and electronics diagrams should be emailed to David and Blake.
    • DAQ/Online/Counting Room: A single TCP/IP port is needed in the Alcove for the IU PC to control the BCAL cart.
  4. Monte Carlo:
    • Simulations done at Regina by Rafael and Blake were discussed. Blake showed plots of the cosmic_muons simulations. These will be used to assist in the gain matching procedure. Note the reduced energy for the Alcove configuration (cosmics enter side of Module). Rafael showed the simulations with all beam elements (target, He bag, etc). See the link titled 'Simulation of the BCAL at Hall-B setup' at [[1]]. A discussion ensued on the percent of e+e- pairs created based on radiation length. Dan Sober promised to calculate this. Rafael will improve his grid plot (Figure 5) to indicate the value of energy deposited in each box.
  5. Any other business
    • Run plan: Regina will work on the Setup, Calibration and Run Plan the week of August 14.
    • Travel Plans: Dan will participate in the September run.
    • 2005 BCAL/TRIUMF Analysis: This is nearing completion and the student (Gergana Koleva) is close to defending her M.Sc.

Action Items

  1. BPM, Veto: David and Elliott will work on making this a functioning system.
  2. Light Leak: Elton will look for black tedlar material (or equivalent).
  3. Cosmics counters: Regina will provide.
  4. Mu metal foil: Eric will measure the field inside the insert and mu-metal by making a single mu-metal shield out of the sheet that Regina has ordered. We will measure the field in the Alcove as a function of BCAL position.
  5. Tagger, Trigger, Cables: The trigger logic and timing details are being continously refined. The fast core cable is pending.
  6. Data storage/Silo Request: David and Zisis will deal with this offline.
  7. Online Macros/MC Data steam/MC Validation & Binning: David and Zisis will work on these offline. Zisis will provide requirements for ROOT trees.
  8. Beam Request: Elton, David and Zisis will carry this discussion offline.
  9. Safety, COOP: We will piggy-back on standard H-B documentation with the addition of our oqn equipment.
  10. Run Plan: Regina will draft this and the calibration procedure with cosmics and beam.